OpenStudy (anonymous):

4/y^2+y-12 = 1/y+4 - 2/y-3

OpenStudy (anonymous):

Please see attachment - is this what you mean?

OpenStudy (anonymous):

i meant 4 over y squared+y- 12 = 1 over (y+4) - 2 over (y-3)

OpenStudy (anonymous):

You mean this? See attachment.

OpenStudy (anonymous):

yes! thats what i meant. how did you get 4= y-2(y+4) -3 by multiplying y^2 + y -1 to both sides?

OpenStudy (anonymous):

Note that \(y^2+y-12=(y+4)(y-3)\)